# Garf: Learning Generalizable 3D Reassembly for Real-World Fractures

### v1

@@ Abstract
3D reassembly is a challenging spatial intelligence task with broad applications across scientific domains. While large-scale synthetic datasets have fueled promising learning-based approaches, their generalizability to different domains is limited. Critically, it remains uncertain whether models trained on synthetic datasets can generalize to real-world fractures where breakage patterns are more complex. To bridge this gap, we propose Garf, a generalizable 3D reassembly framework for real-world fractures. Garf leverages fracture-aware pretraining to learn fracture features from individual fragments, with flow matching enabling precise 6-DoF alignments. At inference time, we introduce one-step preassembly, improving robustness to unseen objects and varying numbers of fractures. In collaboration with archaeologists, paleoanthropologists, and ornithologists, we curate Fractura, a diverse dataset for vision and learning communities, featuring real-world fracture types across ceramics, bones, eggshells, and lithics. Comprehensive experiments have shown our approach consistently outperforms state-of-the-art methods on both synthetic and real-world datasets, achieving 82.87% lower rotation error and 25.15% higher part accuracy. This sheds light on training on synthetic data to advance real-world 3D puzzle solving, demonstrating its strong generalization across unseen object shapes and diverse fracture types. Garf is available here.
@@ 1 Introduction
We have long been captivated by questions of human origins [52]: What are we? Where do we come from? The answers to these fundamental questions lie in archaeological materials such as bones [6], ceramics [50], and lithics [34]. However, these artifacts are often highly fragmented and incomplete [6, 19]. Reassembling them requires placing each fragment in its correct position and orientation to restore a complete or functional entity [41]. This process is not only time-consuming but also challenges the limits of human spatial intelligence [41, 10, 47]. For instance, an experimental study on lithic refitting reported a success rate of only 30%, with experts performing only marginally better than novices [20]. Consequently, museum storerooms around the world remain filled with thousands of unassembled fragments, waiting to be pieced back together [47].
Recently, the emergence of the large-scale dataset Breaking Bad [42] has brought new hope to this domain, fueling the development of data-intensive reassembly methods [54, 49, 61, 23, 30, 7, 41, 51, 28, 57, 21, 17]. While PuzzleFusion++ achieves state-of-the-art (SOTA) performance on the everyday subset, its generalization degrades significantly on the artifact subset due to its reliance on global geometry learning [51]. More critically, the Breaking Bad dataset is generated via physics-based simulation [43], raising a fundamental question: can models trained on synthetic breakage patterns generalize to real-world fractures?
@@
To answer this question, we identify two major real-world fracture challenges and curate Fractura, a dataset capturing key complexities: (i) Data diversity encompasses three geometrically distinct fracture types across multiple scientific domains, including bones, ceramics, eggshells, and lithics. As shown in Fig. 2, ceramics exhibit irregular, chaotic fractures typical of random breakage events, whereas flintknapping produces conchoidal fractures. This allows a systematic study of how object shapes and fracture types affect reassembly performance. (ii) Missing or extraneous fragments are common issues in real-world scenarios [13, 47] (see Fig. 2), which guides our model design to improve robustness.
In response to these complexities, we propose Garf, a generalizable 3D reassembly framework for real-world fractures, featuring four components: (i) Large-scale fracture-aware pretraining takes lessons from recent successes of large-scale pretraining in natural language and computer vision [35, 40, 45, 1, 44]. This module learns fracture segmentation of individual fragments to handle unseen objects as well as missing or extraneous fragments. (ii) Flow-based reassembly on $\mathrm{SE}(3)$ introduces flow matching to learn pose distribution, leveraging the $\mathrm{SO}(3)$ manifold for accurate rotation estimation. Inspired by human puzzle-solving, we design a multi-anchor training strategy that randomly selects a subset of fragments to form local structures, exposing the model to diverse combinations to enhance distribution learning. (iii) One-step pre-assembly at inference time emphasizes the importance of pose initialization. Unlike image generation methods using a verifier to select optimal noise, we observe that the one-step output of flow matching provides a strong pose initialization. (iv) LoRA-based fine-tuning enhances the model’s adaptability to specific domains.
@@
  • We introduce Garf, the first flow-based 3D fracture assembly framework that integrates fracture-aware pretraining and test-time one-step pre-assembly, achieving SOTA performance across synthetic and real-world datasets. Garf reduces rotation error by 82.87% and improves part accuracy by 25.15% compared to previous methods.
@@ 2 Fractura Dataset
Existing datasets, whether synthetic [42] or real [18, 47], are limited to a single fracture type and fail to capture real-world challenges such as missing or extraneous parts. To fill this gap, we curate Fractura, a challenging dataset (see Fig. 2), designed for a comprehensive evaluation of how object shapes, fracture types, and the presence of missing or extraneous parts affect reassembly performance.
Data Characteristics. Fractura comprises both real and synthetic fracture subsets. In collaboration with archaeologists, paleoanthropologists, and ornithologists, the real fracture subset includes three real fracture types relevant to scientific challenges (see Fig. 2): (i) Random breakage produces irregular, chaotic fractures, commonly observed in ceramics, bones, and eggshells. (ii) Incomplete ossification results in unfused bone ends (epiphyses) in juvenile skeletons, leading to fragmented rather than intact bones in skeletal collections and fossil records. Reassembling these unfused parts will benefit the following analysis and further create a complete series of bone developments over time from early childhood to adulthood. (iii) Flintnapping produces conchoidal fractures in lithics, characterized by radially propagating fracture lines (see Fig. 2). Real-world collections and scans naturally incorporate missing or extraneous fragments. For the synthetic fracture subset, we generate realistic fractures using physics-based simulation [43] for ceramics, bones, and eggshells, and geometry-based simulation [36] for lithics.
@@
Datasets Breaking Fantastic RePAIR [47] Fractura
Bad [42] Breaks [18]
@@
Miss. / Extra. $\times$ $\times$ $\times$ $\checkmark$
@@ 3 Method
Previous work either enhances global geometry learning through fragment features [54, 41, 51] or jointly learns hierarchical features from both global and local geometry [30]. In contrast, as shown in Fig. 3, Garf decouples the understanding of local fracture features (Sec. 3.1) and global fragment alignment (Sec. 3.2). At inference, we propose a one-step pre-assembly strategy (Sec. 3.3) for robustness to unseen objects and increasing numbers of fractures. To further boost the performance on domain-specific data, we employ a LoRA-based fine-tuning method (Sec. 3.4).
@@ 3.1 Why Large-Scale Fracture-Aware Pretraining?
Figure 3: Pipeline of Garf. Our framework comprises two main components: (i) Fracture-aware pretraining leverages 14×14$\times$14 × more data than previous methods to learn the local fracture features via fracture point segmentation, and (ii) Flow-based reassembly on $\mathrm{SE}(3)$ leverages the $\mathrm{SO}(3)$ manifold for precise rotation estimation. At inference time, one-step pre-assembly strategy provides better initial poses, enhancing robustness against unseen objects and increasing numbers of fractures.
Humans can infer fracture points without knowledge of the global object shape [37, 30]. To emulate this ability, we leverage large-scale data to learn local fracture features from individual fragments, drawing inspiration from recent advances in large-scale pretraining [35, 40, 45, 1, 44]. This enables our method to generalize to unseen object shapes and handle missing or extraneous fragments.
Specifically, we sample a set of point clouds $\bm{P}=\{P^{1},P^{2},\dots,P^{N}\}$ to represent fragments, where $N$ is the number of fragments. To extract fracture-level features from $\bm{P}$, we employ Point Transformer V3 (PTv3) [55] as the backbone, integrating two MLP layers [30] as the segmentation head for fracture cloud segmentation. Since fracture points constitute only a small proportion of $\bm{P}$, an imbalance arises between positive and negative samples. To mitigate this, we adopt the Dice loss function [46]:
@@
where $p$ is the predicted value and $g$ is its ground truth label. To derive the high-quality $g$, we directly extract shared surfaces between connected fragments from the mesh, defining them as fracture surfaces $\bm{F}$. We then apply Poisson disk sampling [5] to generate $\bm{P}$, ensuring that $g=\bm{P}\cap\bm{F}$. This uniform sampling prevents the encoder from overfitting to specific point densities, improving generalization.
@@ 3.2 Flow-Based Reassembly on SE(3)
Inspired by flow-based generative models [25, 27, 38, 26] in image synthesis [9], protein structure generation [4, 16, 11], and robotic manipulation [3, 58], we leverage flow matching (FM) to model the fracture reassembly process.
@@
The feature $\bm{d}$ is then processed through $L$ Transformer layers including self-attention and global attention. Self-attention adopts efficient execution of FlashAttention [8]:
@@
where the model computes the sequence lengths $\ell$ of the variable number of sampled points per fragment batch, along with their cumulative sums $\operatorname{cu}(\ell)$. The self-attention output $\bm{A}_{\text{self}}$ is added to the original feature $\bm{d}$, yielding the updated representation $\bm{h}=\bm{d}+\bm{A}_{\text{self}}$. The global attention layer then applies a similar attention mechanism to aggregate information across fragments with $\ell=M$. The output undergoes Layer Normalization (LN) and a feed-forward network (FFN) with a residual connection, before regressing the pose $\bm{T}_{t-1}:\{\bm{r}_{t-1},\bm{a}_{t-1}\}$ at the next timestep:
@@
Multi-Anchor Training Strategy. We observe that probability paths vary significantly across fragments; some require complex transformations, while others remain nearly stationary. To model this variation, we introduce a multi-anchor training strategy, randomly selecting $k\in[1,N-1]$ fragments as anchors and fixing their positions with identity rotations and zero translations. For these anchor fragments $\bm{i}$, the vector field is explicitly supervised to be zero: $\bm{v}^{\bm{i}}(\bm{T}_{t},t)=0$. Unlike prior approaches [51], which prevent gradient propagation for a single anchor (the largest fragment and its connected neighbors within a 50% threshold), our multi-anchor strategy enforces a zero vector field, expanding the range of probability paths and enhancing generalization across diverse fragment configurations.
3.3 One-Step Pre-Assembly at Inference Time
In diffusion-based and flow-based models, prior work in image generation emphasizes the role of inference-time scaling, employing search strategies and self-supervised verifiers to find better noise candidates [32, 62, 24, 39]. In contrast, we propose a simple yet effective one-step pre-assembly at inference time, leveraging FM for its ability to model straight-line probability paths. Specifically, we perform a one-step FM inference ($\text{step}=1$) to generate an initial pose $\bm{T}_{0}^{\prime}$, followed by the standard multi-step flow for refinement. This pre-assembly step effectively narrows the search space, improving pose initialization for subsequent refinement. Despite a minimal 5% increase in computational cost ($20\rightarrow 1+20$ steps), our experiments show that this strategy significantly enhances assembly performance, particularly for larger fragment sets.
@@ 3.4 LoRA-based Fine-tuning
To quickly adapt to domain-specific contexts, we employ a LoRA-based [15] fine-tuning approach using the synthetic fracture subset in Fractura. Specifically, we integrate LoRA adapters into the self-attention and global attention layers of the final Transformer block while unfreezing the MLP heads for pose prediction ($f_{\text{rot}}$ and $f_{\text{trans}}$). Our experiments demonstrate that this lightweight fine-tuning method requires as few as 5–10 domain-specific objects to achieve substantial improvements in scientific applications such as juvenile skeleton reconstruction and lithic refitting.
@@ 4 Experiments
Training. For fracture-aware pretraining, Point Transformer V3 (PTv3) [55] serves as the backbone, extracting 64-channel features from its final layer. Garf is pretrained on three Breaking Bad [42] subsets, totaling 1.9M fragments—14×14$\times$14 × more than prior works. For fair comparisons, Garf-mini is pretrained only on the Everyday subset. For FM, both Garf and Garf-mini are trained on the Everyday subset. We use a standard transformer [51] to compute the vector field, with each block consisting of 6 encoder layers, 8 attention heads per layer, and an embedding dimension of 512. The initial learning rate is 2e-4 and decays by a factor of 2 at epochs 900 and 1200. Garf is trained with a batch size of 32 on 4 NVIDIA H100 GPUs, requiring 2 days for pretraining and 3 days for FM. Garf-mini completes pretraining in 0.5 days. For LoRA fine-tuning, we use the PEFT framework [33] with rank $r=128$, $\alpha=256$, and a dropout rate of 0.1.
Datasets. We evaluate our model on three datasets with diverse object shapes and fracture types: (i) Breaking Bad [42], the largest synthetic fracture dataset for 3D reassembly. We use the volume-constrained version, evaluating 7,872 assemblies from the Everyday subset and 3,697 assemblies from the Artifact subset. Results on the vanilla version are in the supplementary materials. (ii) Fantastic Breaks [18], a real-world dataset of 195 manually scanned fractured objects with complex surfaces, used for evaluation only. (iii) Fractura, a mixed synthetic-real dataset spanning ceramics, bones (vertebrae, limbs, ribs), eggshells, and lithics. For the synthetic fracture subset, we follow an 80/20 split [47] for LoRA fine-tuning and evaluation.
Evaluation Metrics. Following [51], we evaluate assembly quality using four metrics: (i) RMSE(R) is the root mean square error of rotation (degrees); (ii) RMSE(T) is the root mean square error of translation; (iii) PA is the percentage of correctly assembled fragments, where the per-fragment chamfer distance is below 0.01; and (iv) CD is the chamfer distance between the assembled object and ground truth.
Competing Methods. We compare our approach against Global [22], LSTM [53], DGL [60], Jigsaw [30], PMTR***The performance of PMTR [21] is from PuzzleFusion++ [51]. [21], and PuzzleFusion++ (PF++)  [51]. We implemented PF++ [51] and Jigsaw [30] using its official codebase, while performance metrics for other methods on the volume-constrained Breaking Bad dataset are sourced from their official papers or repositories. Additional comparisons with SE(3)-Equiv [54], DiffAssemble [41], and PHFormer [7] on the vanilla Breaking Bad dataset, as well as FragmentDiff [57] on its custom Breaking Bad subset, are provided in the supplementary materials.

### v2

@@ Abstract
3D reassembly is a challenging spatial intelligence task with broad applications across scientific domains. While large-scale synthetic datasets have fueled promising learning-based approaches, their generalizability to different domains is limited. Critically, it remains uncertain whether models trained on synthetic datasets can generalize to real-world fractures where breakage patterns are more complex. To bridge this gap, we propose Garf, a generalizable 3D reassembly framework for real-world fractures. Garf leverages fracture-aware pretraining to learn fracture features from individual fragments, with flow matching enabling precise 6-DoF alignments. At inference time, we introduce two-session flow matching, improving robustness to unseen objects and varying numbers of fractures. In collaboration with archaeologists, paleoanthropologists, and ornithologists, we curate Fractura, a diverse dataset for vision and learning communities, featuring real-world fracture types across ceramics, bones, eggshells, and lithics. Comprehensive experiments have shown our approach consistently outperforms state-of-the-art methods on both synthetic and real-world datasets, achieving 82.87% lower rotation error and 25.15% higher part accuracy on the Breaking Bad Everyday dataset. This sheds light on training on synthetic data to advance real-world 3D puzzle solving, demonstrating its strong generalization across unseen object shapes and diverse fracture types. Garf’s code, data and demo are available at https://ai4ce.github.io/GARF/.
@@ 1 Introduction
We have long been captivated by questions of human origins [wilson2012social]: What are we? Where do we come from? The answers to these fundamental questions lie in archaeological materials such as bones [callaway2017oldest], ceramics [wang2019emergence], and lithics [mcpherron2010evidence]. However, these artifacts are often highly fragmented and incomplete [callaway2017oldest, laughlin_149_2005]. Reassembling them requires placing each fragment in its correct position and orientation to restore a complete or functional entity [scarpellini2024diffassemble]. This process is not only time-consuming but also challenges the limits of human spatial intelligence [scarpellini2024diffassemble, gardner2011frames, tsesmelis2025re]. For instance, an experimental study on lithic refitting reported a success rate of only 30%, with experts performing only marginally better than novices [laughlin_experimental_2010]. Consequently, museum storerooms around the world remain filled with thousands of unassembled fragments, waiting to be pieced back together [tsesmelis2025re].
Recently, the emergence of the large-scale dataset Breaking Bad [sellan2022breaking] has brought new hope to this domain, fueling the development of data-intensive reassembly methods [wu2023leveraging, villegas2023matchmakernet, zhang2024scalable, li2024geometric, lu2024jigsaw, cui2024phformer, scarpellini2024diffassemble, wang2024puzzlefusion++, lu2024jigsaw++, xu2024fragmentdiff, lee20243d, kim2024fracture]. While PuzzleFusion++ achieves state-of-the-art (SOTA) performance on the everyday subset, its generalization degrades significantly on the artifact subset due to its reliance on global geometry learning [wang2024puzzlefusion++]. More critically, the Breaking Bad dataset is generated via physics-based simulation [sellan2023breaking], raising a fundamental question: can models trained on synthetic breakage patterns generalize to real-world fractures?
@@
To answer this question, we identify two major real-world fracture challenges and curate Fractura, a dataset capturing key complexities: (i) Data diversity encompasses three geometrically distinct fracture types across multiple scientific domains, including bones, ceramics, eggshells, and lithics. As shown in Fig. 2, ceramics exhibit irregular, chaotic fractures typical of random breakage events, whereas flintknapping produces conchoidal fractures. This allows a systematic study of how object shapes and fracture types affect reassembly performance. (ii) Missing or extraneous fragments are common issues in real-world scenarios [harvati2019apidima, tsesmelis2025re] (see Fig. 2), which guides our model design to improve robustness.
In response to these complexities, we propose Garf, a generalizable 3D reassembly framework for real-world fractures, featuring four components: (i) Large-scale fracture-aware pretraining takes lessons from recent successes of large-scale pretraining in natural language and computer vision [oquab2023dinov2, radford2021learning, stevens2024bioclip, achiam2023gpt, steiner2024paligemma]. This module learns fracture segmentation of individual fragments to handle unseen objects as well as missing or extraneous fragments. (ii) Flow-based reassembly on $\mathrm{SE}(3)$ introduces flow matching to learn pose distribution, leveraging the $\mathrm{SO}(3)$ manifold for accurate rotation estimation. Inspired by human puzzle-solving, we design a multi-anchor training strategy that randomly selects a subset of fragments to form local structures, exposing the model to diverse combinations to enhance distribution learning. (iii) Two-session flow matching at inference time emphasizes the importance of pose initialization. The first session performs a one-step flow matching to generate a strong pose initialization on SE(3); the second session performs standard multi-step flow for refinement. (iv) LoRA-based fine-tuning enhances the model’s adaptability to specific domains.
@@
  • We introduce Garf, the first flow-based 3D fracture assembly framework that integrates fracture-aware pretraining and two-session flow matching, achieving SOTA performance across synthetic and real-world datasets.
@@ 2 Fractura Dataset
Existing datasets, whether synthetic [sellan2022breaking] or real [lamb2023fantastic, tsesmelis2025re], are limited to a single fracture type. To fill this gap, we curate Fractura, a challenging dataset (see Fig. 2), designed for a comprehensive evaluation of how object shapes, fracture types, and the presence of missing or extraneous parts affect reassembly performance.
Data Characteristics. Fractura comprises both real and synthetic fracture subsets. In collaboration with archaeologists, paleoanthropologists, and ornithologists, the real fracture subset includes three real fracture types relevant to scientific challenges (see Fig. 2): (i) Random breakage produces irregular, chaotic fractures, commonly observed in ceramics, bones, and eggshells. (ii) Incomplete ossification results in unfused bone ends (epiphyses) in juvenile skeletons, leading to fragmented rather than intact bones in skeletal collections and fossil records. Reassembling these unfused parts will benefit the following analysis and further create a complete series of bone developments over time from early childhood to adulthood. (iii) Flintnapping produces conchoidal fractures in lithics, characterized by radially propagating fracture lines (see Fig. 2). Real-world collections and scans naturally incorporate missing or extraneous fragments. For the synthetic fracture subset, we generate realistic fractures using physics-based simulation [sellan2023breaking] for ceramics, bones, and eggshells, and geometry-based simulation [orellana2021proof] for lithics.
@@
Datasets Breaking Fantastic RePAIR [tsesmelis2025re] Fractura
Bad [sellan2022breaking] Breaks [lamb2023fantastic]
@@
Miss. / Extra. $\times$ / $\times$ $\times$ / $\times$ $\checkmark$ / $\times$ $\checkmark$ / $\checkmark$
@@ 3 Method
Previous work either enhances global geometry learning through fragment features [wu2023leveraging, scarpellini2024diffassemble, wang2024puzzlefusion++] or jointly learns hierarchical features from both global and local geometry [lu2024jigsaw]. In contrast, as shown in Fig. 3, Garf decouples the understanding of local fracture features (Sec. 3.1) and global fragment alignment (Sec. 3.2). At inference, we propose the two-session flow matching (Sec. 3.3) for robustness to unseen objects and increasing numbers of fractures. To further boost the performance on domain-specific data, we employ a LoRA-based fine-tuning method (Sec. 3.4).
@@ 3.1 Why Large-Scale Fracture-Aware Pretraining?
Figure 3: Pipeline of Garf. Our framework comprises two main components: (i) Fracture-aware pretraining leverages $14\times$ more data than previous methods to learn the local fracture features via fracture point segmentation, and (ii) Two-session flow-based reassembly on $\mathrm{SE}(3)$ leverages the $\mathrm{SO}(3)$ manifold for precise rotation estimation. At inference time, one-step pre-assembly strategy provides better initial poses, enhancing robustness against unseen objects and increasing numbers of fractures.
Humans can infer fracture points without knowledge of the global object shape [papaioannou2001virtual, lu2024jigsaw]. To emulate this ability, we leverage large-scale data to learn local fracture features from individual fragments, drawing inspiration from recent advances in large-scale pretraining [oquab2023dinov2, radford2021learning, stevens2024bioclip, achiam2023gpt, steiner2024paligemma]. This enables our method to generalize to unseen object shapes and handle missing or extraneous fragments.
Specifically, we sample a set of point clouds $\bm{P}=\{P^{1},P^{2},\dots,P^{N}\}$ to represent fragments, where $N$ is the number of fragments. To extract fracture-level features from $\bm{P}$, we employ Point Transformer V3 (PTv3) [wu2024point] as the backbone, integrating two MLP layers [lu2024jigsaw] as the segmentation head for fracture cloud segmentation. Since fracture points constitute only a small proportion of $\bm{P}$, an imbalance arises between positive and negative samples. To mitigate this, we adopt the Dice loss function [sudre2017generalised]:
@@
where $p$ is the predicted value and $g$ is its ground truth label. To derive the high-quality $g$, we directly extract shared surfaces between connected fragments from the mesh, defining them as fracture surfaces $\bm{F}$. We then apply Poisson disk sampling [bridson2007fast] to generate $\bm{P}$, ensuring that $g=\bm{P}\cap\bm{F}$. This weighted sampling based on the surface area of fragments prevents the encoder from overfitting to specific point densities, improving generalization.
@@ 3.2 Flow-Based Reassembly on SE(3)
Inspired by flow-based generative models [lipman2022flow, liu2022flow, pooladian2023multisample, lipman2024flow] in image synthesis [esser2024scaling], protein structure generation [bose2023se, huguet2024sequence, geffnerproteina], and robotic manipulation [black2024pi0, yim2023fast], we leverage flow matching (FM) to model the fracture reassembly process.
@@
The feature $\bm{d}$ is then processed through $L$ Transformer layers including self-attention and global attention. Leveraging the efficient computation of FlashAttention [dao2023flashattention2], we modify the self-attention module to allow more attention on large fragments:
@@
where the model computes the sequence lengths $\ell$ of the variable number of sampled points per fragment batch, along with their cumulative sums $\operatorname{cu}(\ell)$. More details are in the supplementary materials. The self-attention output $\bm{A}_{\text{self}}$ is added to the original feature $\bm{d}$, yielding the updated representation $\bm{h}=\bm{d}+\bm{A}_{\text{self}}$. The global attention layer then applies a similar attention mechanism to aggregate information across fragments with $\ell=M$. The output undergoes Layer Normalization (LN) and a feed-forward network (FFN) with a residual connection, before regressing the pose $\bm{T}_{t-1}:\{\bm{r}_{t-1},\bm{a}_{t-1}\}$ at the next timestep:
@@
Multi-Anchor Training Strategy. We observe that probability paths vary significantly across fragments; some require complex transformations, while others remain nearly stationary. To model this variation, we introduce a multi-anchor training strategy, randomly selecting $k\in[1,N-1]$ fragments as anchors and fixing their positions with identity rotations and zero translations. For these anchor fragments $\bm{i}$, the vector field is explicitly supervised to be zero: $\bm{v}^{\bm{i}}(\bm{T}_{t},t)=0$. Unlike prior approaches [wang2024puzzlefusion++], which prevent gradient propagation for a single anchor (the largest fragment and its connected neighbors within a 50% threshold), our multi-anchor strategy enforces a zero vector field, expanding the range of probability paths and enhancing generalization across diverse fragment configurations.
3.3 Two-Session Flow Matching at Inference Time
Prior diffusion/flow matching work only uses single-session inference. While image generation methods use search strategies or self-supervised verifiers to improve initialization [ma2025inference, zhou2024golden, li2024enhancing, qi2024not], such strategies are less effective for pose estimation. In contrast, we design the two-session flow matching at inference time: the first session performs a one-step FM inference ($\text{step}=1$) to generate an initial pose $\bm{T}_{0}^{\prime}$, leveraging FM’s ability to model straight-line probability paths and narrow the search space; the second session performs standard multi-step flow for refinement. Despite a minimal 5% increase in computational cost ($20\rightarrow 1+20$ steps), this design outperforms single-session inference, particularly for larger fragment sets.
@@ 3.4 LoRA-based Fine-tuning
To quickly adapt to domain-specific contexts, we employ a LoRA-based [hu2022lora] fine-tuning approach using the synthetic fracture subset in Fractura. Specifically, we integrate LoRA adapters into the self-attention and global attention layers of the final Transformer block while unfreezing the MLP heads for pose prediction ($f_{\text{rot}}$ and $f_{\text{trans}}$). Our experiments demonstrate that this lightweight fine-tuning method requires as few as 5–10 domain-specific objects to achieve substantial improvements in scientific applications such as juvenile skeleton reconstruction and lithic refitting.
@@ 4 Experiments
[…]
Datasets. We evaluate our model on three datasets with diverse object shapes and fracture types: (i) Breaking Bad [sellan2022breaking], the largest synthetic fracture dataset for 3D reassembly. We use the volume-constrained version, evaluating 7,872 assemblies from the Everyday subset and 3,697 assemblies from the Artifact subset. Results on the vanilla version are in the supplementary materials. (ii) Fantastic Breaks [lamb2023fantastic], a real-world dataset of 195 manually scanned fractured objects with complex surfaces, used for evaluation only. (iii) Fractura, a mixed synthetic-real dataset spanning ceramics, bones (vertebrae, limbs, ribs), eggshells, and lithics. For the synthetic fracture subset, we follow an 80/20 split [tsesmelis2025re] for LoRA fine-tuning and evaluation.
Evaluation Metrics. Following [wang2024puzzlefusion++], we evaluate assembly quality using four metrics: (i) RMSE(R) is the root mean square error of rotation (degrees); (ii) RMSE(T) is the root mean square error of translation; (iii) PA is the percentage of correctly assembled fragments, where the per-fragment chamfer distance is below 0.01; and (iv) CD is the chamfer distance between the assembled object and ground truth.
Competing Methods. We compare our approach against Global [li2020learning], LSTM [wu2020pq], DGL [zhan2020generative], Jigsaw [lu2024jigsaw], PMTR***The performance of PMTR [lee20243d] is from PuzzleFusion++ [wang2024puzzlefusion++]. [lee20243d], and PuzzleFusion++ (PF++)  [wang2024puzzlefusion++]. We implemented PF++ [wang2024puzzlefusion++] and Jigsaw [lu2024jigsaw] using its official codebase, while performance metrics for other methods on the volume-constrained Breaking Bad dataset are sourced from their official papers or repositories. Additional comparisons with SE(3)-Equiv [wu2023leveraging], DiffAssemble [scarpellini2024diffassemble], and PHFormer [cui2024phformer] on the vanilla Breaking Bad dataset, as well as FragmentDiff [xu2024fragmentdiff] on its custom Breaking Bad subset, are provided in the supplementary materials.